OpenStudy (anonymous):

How to solve square root of (x + 4) + square root of (1-x) = 3?

OpenStudy (tkhunny):

1) Think on the Domain 2) Isolate and Square \(\sqrt{x+4}\) requires \(x \ge -4\) \(\sqrt{1-x}\) requires \(x \le 1\) Thus \(x\in [-4,1]\) That's #1. Here's "Isolate". \(\sqrt{x+4} = 3 - \sqrt{1-x}\) Okay, you do the rest.

OpenStudy (anonymous):

Thanks!